OpenStudy (buggiethebug):

I need help with my physics lab. projectile motion lab on an air table that was tilted. the puck was projected (pushed once) across the table in a projectile manor and dots were made with a spark timer on a piece of paper. the dots were recorded as x,y, and time in tables. We have to find the horizontal velocity and acceleration and graphing it. can someone tell me how to do it? I was thinking of using derivatives to find the v-t but, i need to prove that what i have is a parabola(quadratic) how do i prove it?

OpenStudy (anonymous):

To show that a curve is a parabola, show that the df/dx = kx+k1 ie a straight line. In this case show dy/dx = ax+b using the eq of motion for y(t) and x(t) eliminate t to get y=f(x).